# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v3

@@ 1. Introduction
Our approach is to translate the appearing questions to questions about final segments in the value group $vK$ of $(K,v)$. The latter are connected with the investigation of cuts in ordered abelian groups which has appeared at several points in the literature. For a survey and a list of references, see [8]. The present paper uses, and improves, parts of the unpublished manuscript [10]. Let us describe the background on cuts first.
@@ 2.2. Final segments in ordered abelian groups
If $M$, $M_{1}$ and $M_{2}$ are nonempty subsets of $G$, then we set $M_{1}+M_{2}=\{\alpha+\beta\mid\alpha\in M_{1}\,,\,\beta\in M_{2}\}$ and $M_{1}-M_{2}=\{\alpha-\beta\mid\alpha\in M_{1}\,,\,\beta\in M_{2}\}$. From this, the sets $\alpha+M$, $\alpha-M$ and $M-\alpha$ are obtained via replacing $\alpha$ by $\{\alpha\}$. We set $-M:=0-M$.
@@ 3.1. Overrings of $\mathcal{O}_{v}$ and their ideals
This is a convex subgroup of the value group $vK$ of $(K,v)$. The valuation $w$ associated with $\mathcal{O}$ is (up to equivalence) given by
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
[…]
@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}\,$. The annihilator of the quotient $I_{1}/I_{2}$ with $I_{2}\subseteq I_{1}$ is the $\mathcal{O}_{v}$-ideal $I_{2}:I_{1}$. In general, the equation $I_{1}J=I_{2}$ is not solvable and $J=I_{2}:I_{1}$ will not be a solution, but it will be the largest $\mathcal{O}_{v}$-ideal such that $I_{1}J\subseteq I_{2}\,$. However, in the situation appearing in the papers [3, 4], we have $I_{1}=U$ and $I_{2}=UV$ for $\mathcal{O}_{v}$-ideals $U,V$. Then $J=V$ is a solution of $I_{1}J=I_{2}$ and our task is just to compute the largest solution $I_{2}:I_{1}$, which we accomplish in the next theorem.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subseteq I_{1}\,$. If the annihilator of $I_{1}/I_{2}$ is $\mathcal{M}_{v}\,$, then $I_{1}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
1) Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subseteq I_{1}\,$. Then the annihilator of $I_{1}/I_{2}$ is equal to $\mathcal{M}_{v}$ if and only if $I_{1}$ is principal and $I_{2}=I_{1}\mathcal{M}_{v}\,$.
2) Take $\mathcal{O}_{v}$-ideals $I$ and $J$ with $J\subseteq\mathcal{O}_{v}\,$. Then the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subseteq I_{1}\,$.
@@ 3.6. Application to the computation of annihilators in special cases
Throughout, we assume that $2\leq n\in\mathbb{N}$. For an $\mathcal{O}_{v}$-module $M$ we denote by $\mbox{\rm ann}\,\,M$ the annihilator of $M$.
@@
Take an $\mathcal{O}_{v}$-ideal $I\subseteq\mathcal{O}_{v}$ and set $J=bI^{n-1}$, where $b\in\mathcal{O}_{v}\,$.
@@
2) If $\mathcal{M}_{v}$ annihilates $I/IJ$, then $I/IJ=0$, except if $\mathcal{M}_{v}$ is principal and either $I=J=\mathcal{M}_{v}$ (which holds if and only if $n=2$ and $b\in\mathcal{O}_{v}^{\times}$), or $I=\mathcal{O}_{v}$ and $J=\mathcal{M}_{v}=(b)$ (which holds if and only if $n=1$ and $b\notin\mathcal{O}_{v}^{\times}$). In these exceptional cases, $\mathcal{M}_{v}=\mbox{\rm ann}\,I/IJ$.
@@
Take any valuation ring $\mathcal{O}$ with maximal ideal $\mathcal{M}$. Then
@@
1): We have $a\notin a\mathcal{M}$, hence if $a\in\mathcal{M}$, then $a\mathcal{M}\neq\mathcal{M}$. Since $\mathcal{O}_{v}\subseteq\mathcal{O}$, we have $a\in\mathcal{O}$. Thus if $a\notin\mathcal{M}$, then $a$ is a unit in $\mathcal{O}$, so $a\mathcal{M}=\mathcal{M}$.
2): The value group of $w$ is not discrete, and hence dense, if and only if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. If it is discrete and $\gamma$ is its smallest positive element, then =$\mathcal{M}=\{b\in K\mid wb\geq\gamma\}$ and $\mathcal{M}^{n}=\{c\in K\mid wc\geq n\gamma\}\subsetneq\mathcal{M}$ since $n\gamma>\gamma$. If it is dense, then for every $b\in\mathcal{M}$ there is $c\in K$ such that $0<nwc<wb$, whence $b\in\mathcal{M}^{n}$; therefore, $\mathcal{M}^{n}\subseteq\mathcal{M}\subseteq\mathcal{M}^{n}$ and consequently, $\mathcal{M}^{n}=\mathcal{M}$.
@@
Assume that $a\in\mathcal{O}_{v}\,$, take $\mathcal{O}\in\mathcal{R}$ with maximal ideal $\mathcal{M}$ and denote the valuation associated with $\mathcal{O}$ by $w$.
1) The following holds:
$a\mathcal{M}:(a\mathcal{M})^{n}\>=\>\left\{\begin{array}[]{ll}(a\mathcal{M})^{n-1}&\mbox{if $\mathcal{M}$ is a principal $\mathcal{O}$-ideal, }\\ (a\mathcal{O})^{n-1}=a^{n-1}\mathcal{O}&\mbox{if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal.}\end{array}\right.$
2)  We have that $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=\mathcal{M}_{v}$ if and only if $n=2$, $a\notin\mathcal{M}_{v}$ and $\mathcal{M}=\mathcal{M}_{v}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
It remains to show that $J\mathcal{O}(IJ)^{\diamondsuit}=J^{\diamondsuit}$ is equal to $(a\mathcal{O})^{n-1}$ if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. In this case, $J=(a\mathcal{M})^{n-1}=a^{n-1}\mathcal{M}$, and 00 is the infimum of $w\mathcal{M}$, not contained in $w\mathcal{M}$. Therefore, $wa^{n-1}\notin wa^{n-1}+w\mathcal{M}=wa^{n-1}\mathcal{M}$ is the infimum of $wa^{n-1}\mathcal{M}$. Hence by definition, $J^{\diamondsuit}=\{c\mid wc\geq wa^{n-1}\}=a^{n-1}\mathcal{O}=(a\mathcal{O})^{n-1}$. This completes the proof of part 1).

### v4

@@ 1. Introduction
[…]
@@ 2.2. Final segments in ordered abelian groups
[…]
@@ 3.1. Overrings of $\mathcal{O}_{v}$ and their ideals
[…]
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
(45) $\mathcal{O}(\mathcal{O})\>=\>\mathcal{O}(\mathcal{M})\>=\>\mathcal{O}\;\;\;\mbox{\ \ and\ \ }\;\;\;\mathcal{M}(\mathcal{O})\>=\>\mathcal{M}(\mathcal{M})\>=\>\mathcal{M}\>.$
@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}\,$. The annihilator of the quotient $I_{1}/I_{2}$ with $I_{2}\subsetneq I_{1}$ is the $\mathcal{O}_{v}$-ideal $I_{2}:I_{1}$. In general, the equation $I_{1}J=I_{2}$ is not solvable and $J=I_{2}:I_{1}$ will not be a solution, but it will be the largest $\mathcal{O}_{v}$-ideal such that $I_{1}J\subseteq I_{2}\,$. However, in the situation appearing in the papers [3, 4], we have $I_{1}=U$ and $I_{2}=UV$ for $\mathcal{O}_{v}$-ideals $U,V$. Then $J=V$ is a solution of $I_{1}J=I_{2}$ and our task is just to compute the largest solution $I_{2}:I_{1}$, which we accomplish in the next theorem.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subsetneq I_{1}\,$. If the annihilator of $I_{1}/I_{2}$ is $\mathcal{M}_{v}\,$, then $I_{1}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
1) Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subsetneq I_{1}\,$. Then the annihilator of $I_{1}/I_{2}$ is equal to $\mathcal{M}_{v}$ if and only if $I_{1}$ is principal and $I_{2}=I_{1}\mathcal{M}_{v}\,$.
2) Take $\mathcal{O}_{v}$-ideals $I$ and $J$ with $J\subsetneq\mathcal{O}_{v}\,$. Then the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subsetneq I_{1}\,$.
@@ 3.6. Application to the computation of annihilators in special cases
For an $\mathcal{O}_{v}$-module $M$ we denote by $\mbox{\rm ann}\,\,M$ the annihilator of $M$.
@@
Take an $\mathcal{O}_{v}$-ideal $I\subsetneq\mathcal{O}_{v}$ and $n\in\mathbb{N}$, and set $J=bI^{n-1}$, where $b\in\mathcal{O}_{v}\,$. If $n=1$, assume that $b\in\mathcal{M}_{v}\,$. Then $I/IJ=I/bI^{n}$ is not zero, and the following statements hold.
@@
2) We have that $\mbox{\rm ann}\,I/IJ=\mathcal{M}_{v}$ if and only if $\mathcal{M}_{v}$ is principal and either $I=J=\mathcal{M}_{v}$ (which holds if and only if $n=2$ and $b\in\mathcal{O}_{v}^{\times}$), or $I=\mathcal{O}_{v}$ and $J=\mathcal{M}_{v}=(b)$ (which holds if and only if $n=1$ and $b\notin\mathcal{O}_{v}^{\times}$).
@@
Our assumption that $b\in\mathcal{M}_{v}$ in the case of $n=1$ and that $b\in\mathcal{O}_{v}$ and $I\subsetneq\mathcal{O}_{v}$ in the case of $n>1$ implies that $J\subsetneq\mathcal{O}_{v}\,$. This yields that $I/IJ$ is not zero.
@@
Take $n\geq 2$, any valuation ring $\mathcal{O}$ with maximal ideal $\mathcal{M}$, and $a\in\mathcal{O}$. Then
@@
1): We have $a\notin a\mathcal{M}$, hence if $a\in\mathcal{M}$, then $a\mathcal{M}\neq\mathcal{M}$. If $a\notin\mathcal{M}$, then $a$ is a unit in $\mathcal{O}$, so $a\mathcal{M}=\mathcal{M}$.
[…]
@@
Take $n\geq 2$, $a\in\mathcal{O}_{v}$ and $\mathcal{O}\in\mathcal{R}$ with maximal ideal $\mathcal{M}$. Assume that $(a\mathcal{M})^{n}\neq a\mathcal{M}$. Then the following statements hold.
1) We have that
$\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\>=\>\left\{\begin{array}[]{ll}(a\mathcal{M})^{n-1}&\mbox{if $\mathcal{M}$ is a principal $\mathcal{O}$-ideal, }\\ (a\mathcal{O})^{n-1}=a^{n-1}\mathcal{O}&\mbox{if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal.}\end{array}\right.$
2)  The annihilator is equal to $\mathcal{M}_{v}$ if and only if $n=2$, $a\notin\mathcal{M}_{v}$ and $\mathcal{M}=\mathcal{M}_{v}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
It remains to show that $J\mathcal{O}(IJ)^{\diamondsuit}=J^{\diamondsuit}$ is equal to $(a\mathcal{O})^{n-1}$ if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. Denote the valuation associated with $\mathcal{O}$ by $w$. In the present case, $J=(a\mathcal{M})^{n-1}=a^{n-1}\mathcal{M}$, and 0 is the infimum of $w\mathcal{M}$, not contained in $w\mathcal{M}$. Therefore, $wa^{n-1}\notin wa^{n-1}+w\mathcal{M}=wa^{n-1}\mathcal{M}$ is the infimum of $wa^{n-1}\mathcal{M}$. Hence by definition, $J^{\diamondsuit}=\{c\mid wc\geq wa^{n-1}\}=a^{n-1}\mathcal{O}=(a\mathcal{O})^{n-1}$. This completes the proof of part 1).